OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point? m=-9 (1,3) y=?

myininaya (myininaya):

y=mx+b y=-9x+b we know a point on the line (x,y)=(1,3) plug those in and solve for b

OpenStudy (anonymous):

y-y1=m(x-x1) y-3=(-9)(x-1) y-3=9-x y=12-x or y=-x+12

OpenStudy (anonymous):

y-3=-9(x-1) y = -9x +9+3 y= -9x+12

OpenStudy (anonymous):

Nancy is right... i apparently can't multiply. its -9x, not -x